DIFFERENTIAL GEOMETRY/PDE SEMINAR

Thursday, March 11, 2004 Padelford C-36 1:30 pm

Unbounded bi-invariant quasi-norm on groups of geometric origin

Dmitri Burago (Pennsylvania State University)

By UBSN, unbounded bi-invarian (quasi) semi norm, we mean an unbounded function from a group to reals which satisfies the definition of a bi-invariant semi-norm up to uniformly bounded additive error. We say that a group is unbounded if it admints a UBSN. I will discuss exsamples of bounded and unbounded groups (proving that a specific group is (un)bounded is often quite a non-trivial task), give a number of motivations and applications, and show connections with geometry and sequential dynamics. I will also formulate a number of open questions, which form the most important part of this talk. This is a joint project with S. Ivanow and L. Polterovich.

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.